Prove that 2^n >1+n√[2^(n-1)] for n>2, using the Inequalities
1
Expert's answer
2019-04-03T11:06:18-0400

"2^n >1+n\sqrt{2^{n-1}}" for "n>2"

"2^n-1-n\sqrt{2^{n-1}}>0"

taking derivative of the left part

"2^n*ln(2)-(n*\sqrt{2^{n-1}}*ln(2))/2-\sqrt{2^{n-1}}"

by looking up/making a graph of this function you can see that for n>2 its always positive. It means that that left part will always grow for n>2, so now we only need to check for the inequality to be true for n=2,


"2^2-1-2\sqrt{2^{2-1}}>0"

"3-2*\sqrt2>0"

"\sqrt2" is approximately 1.414 so the inequality is true for n=2 and any n>2